Solve the system of equations using the substitution method.
If the answer is a unique solution, present it as an ordered pair: (x,y). If not, specify whether the answer is "no solution" or "infinitely many solution."
7x+y=1
3x+2y=-9

Answer by nabla(475)

from the first:
y=1-7x
Put that y into the second equation:
3x+2(1-7x)=-9
3x+2-14x=-9
-11x=-11
x=1

If x=1,
y=1-7(1)=-6
check:
7-6=1
3-12=-9
yes.